OpenStudy (anonymous):

What is the equation of the line in slope-intercept form that passes through the point (3, –1) and is perpendicular to the line y = x + 1?

OpenStudy (anonymous):

Do the same as before, only perpendicular lines have negative slopes...so the slope of perpendicular line is -1

OpenStudy (helder_edwin):

actually if u have two lines with slopes m1 and m2 then for them to be perpendicular u gotta have \[ \large m_1\cdot m_2=-1 \]

OpenStudy (anonymous):

yes....so m1 = -1(m2)...sorry if I mis-worded above...this was my intention

OpenStudy (anonymous):

so y=-x+2 is the answer?

OpenStudy (helder_edwin):

yes

OpenStudy (anonymous):

thank you